OpenStudy (anonymous):

OK need more help What is 3/8 +5/7? is it 8/15.

OpenStudy (anonymous):

Ty M

OpenStudy (konradzuse):

That is not correct.

OpenStudy (konradzuse):

both denominators must be the same.

OpenStudy (konradzuse):

when we add and subtract they must be the same, multiplying you do not.

OpenStudy (cwrw238):

you must find the lowest common multiple ( LCM of 8 and 7

OpenStudy (anonymous):

so would it be 1/7?

OpenStudy (cwrw238):

you can only add the top numbers when the denominator is the same

OpenStudy (konradzuse):

nope that isn't correct.

OpenStudy (cwrw238):

LCM = 7*8 = 56

OpenStudy (konradzuse):

Lets find the LCD(Least Common Denominator) HINT: If we cannot find a number lower, we usually multiply the bottom numbers, so in this case 7*8 =?

OpenStudy (konradzuse):

cwrw trying to steal my Q!

OpenStudy (cwrw238):

now you need to convert 3/8 to something over 56

OpenStudy (anonymous):

so 61/56 ? Im lost

OpenStudy (konradzuse):

correct.

OpenStudy (cwrw238):

ok - sorry - you carry on

OpenStudy (konradzuse):

:p

OpenStudy (anonymous):

ty all

OpenStudy (konradzuse):

Do you undertstand how you got that answer though?

OpenStudy (anonymous):

Yes and no? 56yes but no to sure on the 61

OpenStudy (anonymous):

So if you divide 25/6 by 5 is it 5/6?

OpenStudy (anonymous):

ty

OpenStudy (konradzuse):

well actually I wouldn't say that.

OpenStudy (konradzuse):

normally the trick when you have multiple denoms you do this. A/B/C = A *C/B

OpenStudy (konradzuse):

Sorry I said yes, when I was checking my answer it seemed to give me that, and the other one I thought.

OpenStudy (konradzuse):

So 25 / 6 / 5 would be 25 * 5/6

OpenStudy (anonymous):

so would it be 25/30?

OpenStudy (konradzuse):

What is 25*5

OpenStudy (anonymous):

125

OpenStudy (konradzuse):

/6

OpenStudy (anonymous):

so 125/6?

OpenStudy (konradzuse):

yes.

OpenStudy (konradzuse):

http://www.wolframalpha.com/input/?i=%2825%29%2F%286%2F5%29 Depending how you put it into Wolfram you could get 5/6, though I'm not 100% sure why..

OpenStudy (konradzuse):

@jim_thompson5910

OpenStudy (anonymous):

OK ty

jimthompson5910 (jim_thompson5910):

3/8 + 5/7 (7/7)*(3/8) +5/7 21/56 +5/7 21/56 + (5/7)*(8/8) 21/56 + 40/56 (21+40)/56 61/56 So 3/8 + 5/7 = 61/56

OpenStudy (konradzuse):

haha not that part :P. 25/6/5 :)